OpenStudy (anonymous):

Please help me out. :) Find the dimensions of the right circular cone of least volume that can circumscribed about a sphere of radius a. (HINT: let x + 2a be the altitude)

OpenStudy (anonymous):

Let R be the base radius of the cone, h its height and α the half angle at the apex (i.e the angle between the symmetry axis of the cone and its side). Then: h = r(1+sinα)/sinα R = r(1+sinα)/cosα Volume of cone = V = ⅓πr³(1+sinα)³/(cos²α.sinα) You need to take dV/dα, set it = 0 (for a minimum) and solve for α which will give you the values of h and R from the above formulae.

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

is there any easy solution?